OpenStudy (anonymous):

Can somebody please help and explain to me how to check if the solution is extraneous?

OpenStudy (anonymous):

\[\sqrt{2x+6}-\sqrt{x-1}=2\]

OpenStudy (anonymous):

i would add \(\sqrt{x-1}\) to both sides and start with \[\sqrt{2x+6}=\sqrt{x-1}+2\] then square both sides

OpenStudy (anonymous):

then, although your teacher probably wants you to do something else, i would guess at a solution, which should be easy because there are not too many ways to pick an \(x\) make both \(x-1\) and also \(2x+6\) to be perfect squares

OpenStudy (anonymous):

try \(5\) and see that \(2\times 5+6=16\) is a perfect square (it is the square of \(4\)) and also \(5-1=4\) is also a perfect square got it on the first try

OpenStudy (anonymous):

by checking your answer "extraneous" is another name for a solution that doesn't actually work

OpenStudy (shamil98):

a solution that doesn't work with the original equation

OpenStudy (anonymous):

for example \(5\) is a solution to \[\sqrt{2x+6}-\sqrt{x-1}=2\] because \[\sqrt{2\times 5+6}-\sqrt{5-1}=2\\ \sqrt{16}-\sqrt{4}=2\\ 4-2=2\] is true

OpenStudy (anonymous):

Okay thank you sooo much!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I am public schooled!

OpenStudy (anonymous):

oh ok lol